# On von Neumann regularity of ample groupoid algebras

### v2

@@ 5.2. Algebraic partial skew group rings induced by topological partial actions
Let $X$ be a locally compact Hausdorff totally disconnected space, let $G$ be a discrete group with identity element $e$, and let $K$ be a field. Let $\phi=(\{X_{g}\}_{g\in G},\{\phi_{g}\}_{g\in G})$ be a topological partial action of $G$ on $X$. In particular, each $X_{g}$ is a clopen subset of $X$, and $\phi_{g}\colon X_{g^{-1}}\to X_{g}$ is a homeomorphism with inverse $\phi_{g}^{-1}=\phi_{g^{-1}}$. For any $g\in G$, let $C_{c}(X_{g},K)$ denote the ring of all locally constant compactly supported functions $f\colon X_{g}\to K$. Since $X_{g}$ is open in $X$, we may view $C_{c}(X_{g},K)$ as an ideal in $C_{c}(X,K)$. Let $D_{g}=C_{c}(X_{g},K)$ and define $\varphi_{g}\colon D_{g^{-1}}\to D_{g}$ by $\varphi_{g}(f)(x)=f\circ\phi_{g^{-1}}(x)$. Then $\varphi_{g}$ is an isomorphism of $K$-algebras and $\varphi=(\{D_{g}\}_{g\in G},\{\varphi_{g}\}_{g\in G})$ is an algebraic partial action of $G$ on $C_{c}(X,K)$. The partial skew group ring is the $K$-algebra with underlying $K$-vector space
@@
Let $\mathscr{G}$ denote the transformation groupoid associated with $\phi$, and let $e\in G$ denote the identity element. By Theorem 4.2, $C_{c}(X,K)\rtimes_{\varphi}G\cong K\mathscr{G}$ is graded regular if and only if $K\mathscr{G}_{e}\cong C_{c}(X,K)$ is regular. But $C_{c}(X,K)$ is trivially regular since if $f\in C_{c}(X,K)$ and then $ff^{\prime}f=f$ where $f^{\prime}(x)=1/f(x)$ whenever $f(x)\neq 0$, and $f^{\prime}(x)=0$ otherwise. ∎
@@ 5.3.3. Partial action on the tight filters
Suppose $L_{K}(\mathscr{E})\cong K\mathscr{G}$ is regular. Then the isotropy subgroups of $\mathscr{G}$ are locally finite by Theorem 3.1. Since all isotropy subgroups of $\mathscr{G}$ are infinite cyclic, it follows that $\mathscr{G}$ is principal. ∎
@@
Let $\mathscr{E}=(E,\mathscr{L},\mathscr{B})$ be normal labeled space and $K$ a field. Suppose that if $\xi^{\alpha}\in\mathsf{T}$ and $\alpha\neq\beta\gamma^{\infty}$ for any $\beta,\gamma\in\mathscr{L}^{*}$. Then $L_{K}(\mathscr{E})$ is regular.
@@
Notice that the assumption $\alpha\neq\beta\gamma^{\infty}$ for any $\beta,\gamma\in\mathscr{L}^{*}$ implies that all isotropy in $\mathscr{G}_{\mathscr{E}}$ are trivial [14, Proposition 6.10]. Thus, the order of no isotropy element is divisible by the characteristic of $K$. Hence, $L_{K}(\mathscr{E})\cong K\mathscr{G}_{\mathscr{E}}$ is regular by Theorem 3.12. ∎
@@ 5.4. Deaconu-Renault groupoids
The action is said to be directed if, for all $p,q\in P$ with $U_{p}\cap U_{q}\neq\emptyset$, there is $r\in pP\cap qP$ with $U_{r}=U_{p}\cap U_{q}$. Notice that left reversibility is equivalent to this condition if $U_{p}=X$ for all $p\in P$. More generally, if $F\subseteq P$, then $F$ is said to be action directed if given $p,q\in F$ with $U_{p}\cap U_{q}\neq\emptyset$, there is $r\in F\cap pP\cap qP$ with $U_{p}\cap U_{q}=U_{r}$. It is shown in [30, Claim 3] that every finite subset of $P$ is contained in an action-directed finite subset of $P$.
@@
A basis for the topology on $\mathscr{G}(X,P)$ consists of all sets of the form $D(U,p,q,V)=\{(x,pq^{-1},y)\mid x\in U,y\in V,xp=yq\}$ with $U$ and $V$ open subsets of $X$. Such a set is a compact open bisection if $U\subseteq U_{p}$, $V\subseteq V_{q}$ are compact open, $p$ is injective on $U$, $q$ is injective on $V$ and $Up=Vq$. These compact open bisections form a basis for the topology, and so the groupoid is ample. The topology on $\mathscr{G}(X,P)$ is finer than the product topology on $X\times G\times X$ but coincides with the product topology on sets of the form $D(U,p,q,V)$ (see [30, Lemma 5.11]). The groupoid $\mathscr{G}(X,P)$ is Hausdorff. There is a continuous cocycle $c\colon\mathscr{G}(X,P)\to G$ given by $c(x,g,y)=g$ (see [30, Proposition 5.12]).
@@
Let $K$ be a field and let $\mathscr{G}(X,P)$ be the Deaconu-Renault groupoid of the directed action of a cancellative left reversible monoid $P$ on a locally compact, Hausdorff and totally disconnected space $X$ by partial local homeomorphisms. Assume that the domain $U_{p}$ of the action of $p$ is clopen for all $p\in P$. Then $K\mathscr{G}(X,P)$ is graded regular with respect to the $G$-grading coming from the canonical cocycle, where $G$ is the group of fractions of $P$.
@@
Note that $c^{-1}(e)=\{(x,e,y)\in\mathscr{G}(X,P)\}$ and hence is a principal groupoid. So it suffices by Theorem 4.3 to show that $c^{-1}(e)$ is approximately quasi-compact. Let $W$ be a compact open subset of $X$, and let $F$ be an action-directed finite subset of $P$. Put $\mathscr{G}(W,F)=\{(x,e,y)\mid x,y\in W,\exists p\in F,xp=yp\}$. We claim that $\mathscr{G}(W,F)$ is a compact open subgroupoid of $c^{-1}(e)$. First of all,
@@
and hence is compact open, since $W\cap U_{p}$ is compact open because $U_{p}$ is clopen, and the topology on $D(W\cap U_{p},p,p,W\cap U_{p})$ is the product topology on $(W\cap U_{p})\times\{e\}\times(W\cap U_{p})$ by [30, Lemma 5.11]. Indeed, since $X$ is Hausdorff, the set of elements of $U_{p}\times U_{p}$ with $xp=yp$ is closed in $U_{p}\times U_{p}$ and hence a compact subset of $(W\cap U_{p})\times(W\cap U_{p})$ as $U_{p}$ is clopen. It remains to check that $\mathscr{G}(W,F)$ is a subgroupoid. It is clearly closed under inversion. But if $(x,e,y),(y,e,z)\in\mathscr{G}(W,F)$, then we can find $p,q\in F$ with $xp=yp$ and $yq=zq$. Since $F$ is action directed, we can find $r\in F\cap pP\cap qP$ with $U_{r}=U_{p}\cap U_{q}$. Then $y\in U_{r}$. Let $r=pa=qb$. Then $yr$ defined implies that $a$ is defined on $yp=xp$ and $b$ is defined on $yq=zq$. Therefore, $xr=xpa$ and $zr=zqb$ are defined and $xr=xpa=ypa=yr=yqb=zqb=zr$. Thus $(x,e,z)\in\mathscr{G}(W,F)$. We conclude that $\mathscr{G}(W,F)$ is a compact open subgroupoid. If $F,F^{\prime}$ are action-directed finite subsets of $P$ and $W,W^{\prime}$ are compact open subsets of $X$, then there is an action-directed finite subset $F^{\prime\prime}$ with $F\cup F^{\prime}\subseteq F^{\prime\prime}$ by [30, Claim 3]. Then $\mathscr{G}(W,F),\mathscr{G}(W^{\prime},F^{\prime})\subseteq\mathscr{G}(W\cup W^{\prime},F^{\prime\prime})$ by construction. If $(x,e,y)\in c^{-1}(e)$, then we can find a compact open set $W$ containing $x,y$ and we can find $p\in P$ with $xp=yp$. Clearly, $\{p\}$ is action directed. It follows that $c^{-1}(e)$ is the directed union of the compact open subgroupoids $\mathscr{G}(W,F)$ and hence is approximately quasi-compact. The result follows. ∎
@@ 5.5. Self-similar group actions on graphs and Nekrashevych-Exel-Pardo algebras
Let $(G,E)$ be a faithful finite state self-similar group action of a group $G$ on a finite directed graph $E$ with no sinks and let $K$ be a field of characteristic 00. Then the Exel-Pardo algebra $K\mathscr{G}_{G,E}$ is graded regular with respect to its usual $\mathbb{Z}$-grading if and only if $G$ is locally finite.
@@
Suppose first that $K\mathscr{G}_{G,E}$ is graded regular. There is a homomorphism $\psi\colon G\to\Gamma_{c}((\mathscr{G}_{G,E})_{0})$ given by $\psi(g)=\{[g(v)gv^{*},\nu]\mid v\in Z(v),v\in E^{0}\}$. Moreover, $\psi$ is injective because the action is faithful, and so if $g\neq 1$, then $g(x)\neq x$ for some $x\in\partial E$, whence $\mathop{\boldsymbol{r}}([g(s(x))gs(x)^{*},x])=gx\neq x=\mathop{\boldsymbol{s}}([g(s(x))gs(x)^{*},x])$. But $\psi(1)=G^{(0)}$. So $\psi(g)\neq\psi(1)$. By Theorem 3.7, $\Gamma_{c}((\mathscr{G}_{G,E})_{0})$ is locally finite and hence $G\cong\psi(G)$ is locally finite.
@@
Acknowledgments
We thank Gilles de Castro for showing Example 5.12 to us.

### v3

@@ 5.2. Algebraic partial skew group rings induced by topological partial actions
Let $X$ be a locally compact Hausdorff totally disconnected space, let $G$ be a discrete group with identity element $e$, and let $K$ be a commutative ring. Let $\phi=(\{X_{g}\}_{g\in G},\{\phi_{g}\}_{g\in G})$ be a topological partial action of $G$ on $X$. In particular, each $X_{g}$ is a clopen subset of $X$, and $\phi_{g}\colon X_{g^{-1}}\to X_{g}$ is a homeomorphism with inverse $\phi_{g}^{-1}=\phi_{g^{-1}}$. For any $g\in G$, let $C_{c}(X_{g},K)$ denote the ring of all locally constant compactly supported functions $f\colon X_{g}\to K$. Since $X_{g}$ is open in $X$, we may view $C_{c}(X_{g},K)$ as an ideal in $C_{c}(X,K)$. Let $D_{g}=C_{c}(X_{g},K)$ and define $\varphi_{g}\colon D_{g^{-1}}\to D_{g}$ by $\varphi_{g}(f)(x)=f\circ\phi_{g^{-1}}(x)$. Then $\varphi_{g}$ is an isomorphism of $K$-algebras and $\varphi=(\{D_{g}\}_{g\in G},\{\varphi_{g}\}_{g\in G})$ is an algebraic partial action of $G$ on $C_{c}(X,K)$. The partial skew group ring is the $K$-algebra with underlying $K$-vector space
@@
Let $\mathscr{G}$ denote the transformation groupoid associated with $\phi$, and let $e\in G$ denote the identity element. By Theorem 4.2, $C_{c}(X,K)\rtimes_{\varphi}G\cong K\mathscr{G}$ is graded regular if and only if $K\mathscr{G}_{e}\cong C_{c}(X,K)$ is regular. But $C_{c}(X,K)$ is trivially regular. Indeed, if $f\in C_{c}(X,K)$, then we can write $f=\sum_{i=1}^{n}k_{i}1_{U_{i}}$ with the $U_{i}$ disjoint compact open sets (as $X$ is Hausdorff). If we choose $k_{i}^{\prime}\in K$ with $k_{i}k^{\prime}_{i}k_{i}=k_{i}$ and put $f^{\prime}=\sum_{i=1}^{n}k_{i}^{\prime}1_{U_{i}}$, then $f^{\prime}\in C_{c}(X,K)$ and $ff^{\prime}f=f$. ∎
@@ 5.3.3. Partial action on the tight filters
Suppose $L_{K}(\mathscr{E})\cong K\mathscr{G}$ is regular. Then the isotropy subgroups of $\mathscr{G}$ are locally finite by Theorem A. Since all isotropy subgroups of $\mathscr{G}$ are infinite cyclic, it follows that $\mathscr{G}$ is principal. ∎
@@
Let $\mathscr{E}=(E,\mathscr{L},\mathscr{B})$ be normal labeled space and $K$ a field. Suppose that if $\xi^{\alpha}\in\mathsf{T}$ and $\alpha\neq\beta\gamma^{\infty}$ for any $\beta,\gamma\in\mathscr{L}^{\geq 1}$. Then $L_{K}(\mathscr{E})$ is regular.
@@
Notice that the assumption $\alpha\neq\beta\gamma^{\infty}$ for any $\beta,\gamma\in\mathscr{L}^{\geq 1}$ implies that all isotropy in $\mathscr{G}_{\mathscr{E}}$ are trivial [14, Proposition 6.10]. Thus, the order of no isotropy element is divisible by the characteristic of $K$. Hence, $L_{K}(\mathscr{E})\cong K\mathscr{G}_{\mathscr{E}}$ is regular by Theorem A. ∎
@@ 5.4. Deaconu-Renault groupoids
The action is said to be directed if, for all $p,q\in P$ with $U_{p}\cap U_{q}\neq\emptyset$, there is $r\in pP\cap qP$ with $U_{r}=U_{p}\cap U_{q}$. Notice that left reversibility is equivalent to this condition if $U_{p}=X$ for all $p\in P$. More generally, if $F\subseteq P$, then $F$ is said to be action directed if given $p,q\in F$ with $U_{p}\cap U_{q}\neq\emptyset$, there is $r\in F\cap pP\cap qP$ with $U_{p}\cap U_{q}=U_{r}$. It is shown in [31, Claim 3] that every finite subset of $P$ is contained in an action-directed finite subset of $P$.
@@
A basis for the topology on $\mathscr{G}(X,P)$ consists of all sets of the form $D(U,p,q,V)=\{(x,pq^{-1},y)\mid x\in U,y\in V,xp=yq\}$ with $U$ and $V$ open subsets of $X$. Such a set is a compact open bisection if $U\subseteq U_{p}$, $V\subseteq V_{q}$ are compact open, $p$ is injective on $U$, $q$ is injective on $V$ and $Up=Vq$. These compact open bisections form a basis for the topology, and so the groupoid is ample. The topology on $\mathscr{G}(X,P)$ is finer than the product topology on $X\times G\times X$ but coincides with the product topology on sets of the form $D(U,p,q,V)$ (see [31, Lemma 5.11]). The groupoid $\mathscr{G}(X,P)$ is Hausdorff. There is a continuous cocycle $c\colon\mathscr{G}(X,P)\to G$ given by $c(x,g,y)=g$ (see [31, Proposition 5.12]).
@@
Let $R$ be a regular unital ring and let $\mathscr{G}(X,P)$ be the Deaconu-Renault groupoid of the directed action of a cancellative left reversible monoid $P$ on a locally compact, Hausdorff and totally disconnected space $X$ by partial local homeomorphisms. Assume that the domain $U_{p}$ of the action of $p$ is clopen for all $p\in P$. Then $R\mathscr{G}(X,P)$ is graded regular with respect to the $G$-grading coming from the canonical cocycle, where $G$ is the group of fractions of $P$.
@@
Note that $c^{-1}(e)=\{(x,e,y)\in\mathscr{G}(X,P)\}$, and hence it is a principal groupoid. So it suffices by Theorem 4.3 to show that $c^{-1}(e)$ is approximately quasi-compact. Let $W$ be a compact open subset of $X$, and let $F$ be an action-directed finite subset of $P$. Put $\mathscr{G}(W,F)=\{(x,e,y)\mid x,y\in W,\exists p\in F,xp=yp\}$. We claim that $\mathscr{G}(W,F)$ is a compact open subgroupoid of $c^{-1}(e)$. First of all,
@@
and hence is compact open, since $W\cap U_{p}$ is compact open because $U_{p}$ is clopen, and the topology on $D(W\cap U_{p},p,p,W\cap U_{p})$ is the product topology on $(W\cap U_{p})\times\{e\}\times(W\cap U_{p})$ by [31, Lemma 5.11]. Indeed, since $X$ is Hausdorff, the set of elements of $U_{p}\times U_{p}$ with $xp=yp$ is closed in $U_{p}\times U_{p}$ and hence a compact subset of $(W\cap U_{p})\times(W\cap U_{p})$ as $U_{p}$ is clopen. It remains to check that $\mathscr{G}(W,F)$ is a subgroupoid. It is clearly closed under inversion. But if $(x,e,y),(y,e,z)\in\mathscr{G}(W,F)$, then we can find $p,q\in F$ with $xp=yp$ and $yq=zq$. Since $F$ is action directed, we can find $r\in F\cap pP\cap qP$ with $U_{r}=U_{p}\cap U_{q}$. Then $y\in U_{r}$. Let $r=pa=qb$. Then $yr$ defined implies that $a$ is defined on $yp=xp$ and $b$ is defined on $yq=zq$. Therefore, $xr=xpa$ and $zr=zqb$ are defined and $xr=xpa=ypa=yr=yqb=zqb=zr$. Thus $(x,e,z)\in\mathscr{G}(W,F)$. We conclude that $\mathscr{G}(W,F)$ is a compact open subgroupoid. If $F,F^{\prime}$ are action-directed finite subsets of $P$ and $W,W^{\prime}$ are compact open subsets of $X$, then there is an action-directed finite subset $F^{\prime\prime}$ with $F\cup F^{\prime}\subseteq F^{\prime\prime}$ by [31, Claim 3]. Then $\mathscr{G}(W,F),\mathscr{G}(W^{\prime},F^{\prime})\subseteq\mathscr{G}(W\cup W^{\prime},F^{\prime\prime})$ by construction. If $(x,e,y)\in c^{-1}(e)$, then we can find a compact open set $W$ containing $x,y$ and we can find $p\in P$ with $xp=yp$. Clearly, $\{p\}$ is action directed. It follows that $c^{-1}(e)$ is the directed union of the compact open subgroupoids $\mathscr{G}(W,F)$ and hence is approximately quasi-compact. The result follows. ∎
@@ 5.5. Self-similar group actions on graphs and Nekrashevych-Exel-Pardo algebras
Let $(G,E)$ be a faithful finite state self-similar group action of a group $G$ on a finite directed graph $E$ with no sinks and let $R$ be a unital ring. Then the Exel-Pardo algebra $K\mathscr{G}_{G,E}$ is graded regular with respect to its usual $\mathbb{Z}$-grading if and only if $G$ is locally finite.
@@
Suppose first that $K\mathscr{G}_{G,E}$ is graded regular. There is a homomorphism $\psi\colon G\to\Gamma_{c}((\mathscr{G}_{G,E})_{0})$ given by $\psi(g)=\{[g(v)gv^{*},\nu]\mid v\in Z(v),v\in E^{0}\}$. Moreover, $\psi$ is injective because the action is faithful, and so if $g\neq 1$, then $g(x)\neq x$ for some $x\in\partial E$, whence $\mathop{\boldsymbol{r}}\nolimits([g(s(x))gs(x)^{*},x])=gx\neq x=\mathop{\boldsymbol{s}}([g(s(x))gs(x)^{*},x])$. But $\psi(1)=G^{(0)}$. So $\psi(g)\neq\psi(1)$. By Theorem A, $\Gamma_{c}((\mathscr{G}_{G,E})_{0})$ is locally finite and hence $G\cong\psi(G)$ is locally finite.
@@
Appendix: A groupoid-free proof of Theorem B
In this appendix we give a groupoid-free proof of Theorem B.
Let $S$ be an inverse semigroup, and let $R$ be a unital ring. Then $S$ acts on the left of $E(S)$ via the rule $s\mathrel{\diamond}e=ses^{*}$. The $\mathscr{R}$-class of $e\in E(S)$ is the set $R_{e}=\{s\in S\mid ss^{*}=e\}$. Dually the $\mathscr{L}$-class of $e$ is $L_{e}=\{s\in S\mid s^{*}s=e\}$. The maximal subgroup of $S$ at an idempotent $e$ is the group of units $G_{e}=R_{e}\cap L_{e}$ of the inverse monoid $eSe$ (with identity $e$). Note that $G_{e}$ acts freely on the left of $R_{e}$ by multiplication since if $g\in G_{e}$ and $s\in R_{e}$, then we have $gss^{*}g^{*}=geg^{*}=e$, and $s=gs$ implies $e=ss^{*}=gss^{*}=ge=g$.
Lemma 5.16.
Let $S$ be an semigroup and suppose that there is an integer $N>0$ with $|L_{e}|\leq N$ for all $e\in E(S)$. Then $S$ is locally finite.
Proof.
The semigroup $S$ acts on the left of $L_{e}$ by partial injective mappings via the left Schützenberger representation:
$s\cdot t=\begin{cases}st,&\text{if}\ st\in L_{e}\\ \text{undefined},&\text{else}\end{cases}$
for $s\in S$ and $t\in L_{e}$. Moreover, the Schützenberger representations separate points [11]. That is, if $s_{1},s_{2}\in S$ act the same in each Schützenberger representation, then $s_{1}=s_{1}\cdot s_{1}^{*}s_{1}=s_{2}\cdot s_{1}^{*}s_{1}$, and so $s_{1}\leq s_{2}$. Dually, we have $s_{2}\leq s_{1}$, and so they are equal. It follows that the points of $S$ can be separated by homomorphisms into the symmetric inverse monoid $I_{N}$ of all partial bijections on an $N$-element set. Thus $S$ is locally finite by  2.2. ∎
Proof of Theorem B.
The sufficiency of these conditions was established in [34] and was proved in Proposition 2.4.
Suppose next that $RS$ is regular. The augmentation $\varepsilon\colon RS\to R$ given by on the $R$-basis $S$ by $s\mapsto 1$ is a surjective ring homomorphism, and hence $R$ is regular (cf. [19, Lemma 1.3]). Next observe that if $M=S^{1}$ is the inverse monoid resulting from adjoining an identity to $S$, then $RS$ is an ideal of $RM$ and we have that $R\cong RM/RS$. Therefore $RM$ is regular by [19, Lemma 1.3] as $R$ and $RS$ are regular. Thus we may assume without loss of generality that $S$ is a monoid.
Suppose that $X$ is a finite subset of $S$ closed under inversion $x\mapsto x^{*}$. We show that $T=\langle X\rangle$ is finite and the order of each maximal subgroup of $T$ is invertible in $R$. Since every element of a maximal subgroup of $S$ belongs to some finitely generated inverse subsemigroup, it will follow that the order of each element of a maximal subgroup of $S$ is invertible in $R$.
There is a left module homomorphism $\pi\colon RS\to RE(S)$ given by $\pi(s)=ss^{*}$ for $s\in S$. Indeed, $\pi(ts)=tss^{*}t^{*}=t\mathrel{\diamond}(ss^{*})=t\mathrel{\diamond}\pi(s)$ for $s,t\in S$. Evidently, $\ker\pi$ is spanned over $R$ by the elements of the form $s-ss^{*}$ with $s\in S$. Let $L\subseteq\ker\pi$ be the left ideal generated by the elements $t-tt^{*}$ with $t\in T$. We claim that $L$ is generated by the finite set $Y=\{x-xx^{*}\mid x\in X\}$. Indeed, we prove by induction on word length that if $t\in T$, then $t-tt^{*}\in RS\cdot Y$. There is nothing to prove if $t\in X$. Assume that $t=t_{0}x$ with $x\in X$ and $t_{0}$ having shorter word length. Since $t_{0}^{*}$ has the same word length as $t_{0}$, because $X$ is closed under inversion, it follows by induction that $x-xx^{*}$ and $t_{0}^{*}-t_{0}^{*}t_{0}$ are in $RS\cdot Y$. Therefore,
$t-tt^{*}=t_{0}(x-xx^{*})-t_{0}xx^{*}(t_{0}^{*}-t_{0}^{*}t_{0})\in RS\cdot Y$
as required. Since $RS$ is regular, it follows from [19, Theorem 1.1] that there is an idempotent $e\in RS$ with $L=RSe$. Let $f=1-e$ and note that $Lf=RSe(1-e)=0$. Therefore, $\pi(f)=\pi(1)-\pi(e)=1$ (as $e\in L\subseteq\ker\pi$), and also $tf=tt^{*}f$ for all $t\in T$ (as $(t-tt^{*})f=0$). Let $f=\sum_{i=1}^{N}r_{i}s_{i}$ with $r_{i}\neq 0$ for all $i$. We claim that $|L_{x}\cap T|\leq N$ for all $x\in E(T)$. It will then follow that $T$ is finite by Lemma 5.16.
Define [^3] Note that $\widehat{f}$ is essentially the image of $f$ under the isomorphism of $RS$ with $R\mathscr{G}_{S}$, cf. [33] $\widehat{f}\colon S\to R$ by letting
$\widehat{f}(s)=\sum_{s\leq s_{i}}r_{i}.$
Notice that since $s\leq s_{i}$ if and only if $s=ss^{*}s_{i}$, if and only if $s=s_{i}s^{*}s$, we deduce that $\widehat{f}(s)$ is coefficient of $s$ in $ss^{*}f$ and also the coefficient of $s$ in $fs^{*}s$. If $x\in E(S)$, then there are at most $N$ elements in the support of $fx$, and so $\widehat{f}$ has at most $N$ elements of $L_{x}$ in its support.
We claim that if $t\in T\cap L_{ss^{*}}$, then $\widehat{f}(s)=\widehat{f}(ts)$. Indeed, $\widehat{f}(s)$ is the coefficient of $s$ in $ss^{*}f$ and $\widehat{f}(ts)$ is the coefficients of $ts$ in $tss^{*}t^{*}f=tss^{*}t^{*}tf=tss^{*}f$ because $t^{*}\in T$ implies $t^{*}f=t^{*}tf$. But since $t^{*}t=ss^{*}$, left multiplication by $t$ is bijective mapping $ss^{*}S=t^{*}tS\to tt^{*}S=tS$ (with inverse left multiplication by $t^{*}$). Therefore, the coefficient of $ts$ in $tss^{*}f$ is the same as the coefficient of $s$ in $ss^{*}f$, and so $\widehat{f}(ts)=\widehat{f}(s)$.
We may now complete the proof. Let $x\in E(T)$. From
$\sum_{i=1}^{N}r_{i}xs_{i}s_{i}^{*}x^{*}=\pi(xf)=x\mathrel{\diamond}\pi(f)=x\pi(f)x^{*}=x1x^{*}=x$ (5.5)
we deduce that there is $s$ in the support of $xf$ with $ss^{*}=x$. Therefore, $\widehat{f}(s)\neq 0$. Now if $t\in L_{x}\cap T$, then $\widehat{f}(ts)=\widehat{f}(s)\neq 0$ by the above. Since if $t\in L_{x}$, we have that $(ts)^{*}(ts)=s^{*}t^{*}ts=s^{*}ss^{*}s=s^{*}s$, all the elements $ts$ with $t\in L_{x}\cap T$ are in the support of $\widehat{f}$, and so there are at most $N$ such elements. On the other hand, if $t_{1},t_{2}\in L_{x}$ and $t_{1}s=t_{2}s$, then $t_{1}=t_{1}x=t_{1}ss^{*}=t_{2}ss^{*}=t_{2}x=t_{2}$. It follows that $L_{x}\cap T$ has at most $N$ elements.
Finally, we show that if $H_{x}=G_{x}\cap T\subseteq L_{x}\cap T$ is the maximal subgroup of $T$ at $x$, then $|H_{x}|$ is invertible in $R$. Indeed, $H_{x}$ acts freely on the left of $R_{x}$. Let $A$ be a transversal for $H_{x}\backslash R_{x}$. Note that (5.5) implies that
$\sum_{s\in R_{x}}\widehat{f}(s)=\sum_{xs_{i}s_{i}^{*}x^{*}=x}r_{i}=1.$
On the other hand, if $s\in R_{x}$, we already observed that $\widehat{f}(ts)=\widehat{f}(s)$ for all $t\in L_{x}\cap T\supseteq H_{x}$. Since $H_{x}$ acts freely on the left of $R_{x}$, it follows that
$1=\sum_{s\in R_{x}}\widehat{f}(s)=\sum_{a\in A}\sum_{h\in H_{x}}f(ha)=|H_{x}|\sum_{a\in A}f(a),$
and so $|H_{x}|$ is invertible in $R$. This completes the proof. ∎